# Understanding and Mitigating Membership Inference Risks of Neural Ordinary Differential Equations

### v1

@@ 2.2. Membership Inference Attacks
  • Yeom et al. (Yeom et al., 2018) formulate an attack which focuses on differences in loss values: for a specific example $z=(x,y)$, a model will be more accurate in its predictions when it has seen $z$ during training. The attack predicts $z$ as a member if and only if, for a threshold $\tau$, the model’s loss on $z$ is below $\tau$; otherwise, it is classified as a non-member.
  • Shokri et al. (Shokri et al., 2017) attack based on shadow models which better approximate the differences in a model’s responses to members versus non-members. When constructing shadow models, the attacker artificially generates datasets for training and testing these models so that the members and non-members are known in advance. The attacker collects responses from shadow models for both members and non-members and trains a classifier to predict membership of $z$ based on the target model’s response.
  • Song and Mittal (Song and Mittal, 2021) leverage the prediction correctness to compute the threshold $\tau$ for identifying membership. Instead of training classifiers to perform attacks, $\tau$ is designed such that correct predictions with high confidence yield the lowest values while confident but incorrect predictions achieve the highest.
  • Watson et al. (Watson et al., 2022) proposed per-example difficulty calibration in which an attacker leverages shadow models trained without a particular example to compute the average confidence level of a model. The average is then subtracted from the example’s confidence obtained from the target model to calibrate the difficulty of the sample.
  • Carlini et al. (Carlini et al., 2022) introduced the Likelihood Ratio Attack (LiRA), which carefully controls how the adversary splits shadow training data. For each sample in the shadow training data, LiRA trains two sets of shadow models: models with the sample included (“in”) and the others without (“out”). The attacker then collects logits from both members and non-members across these shadow models, computes the threshold (or LiRA membership score) as the ratio of the likelihoods of observing the logit for the sample $z$ in the “in” versus “out” models, and uses this score to identify membership.
Metrics. Initial work (Yeom et al., 2018) on membership inference uses inference accuracy as a metric for evaluating attack success. Subsequent works also uses the area under the ROC curve (AUC) as an additional metric to report the balance between true-positive (TPR) and false-positive rates (FPR). Recent work by Carlini (Carlini et al., 2022) proposes evaluating attack success in a worst-case low-FPR regime. Our work employs all these metrics in evaluation.
2.3. Defenses against Membership Inference
@@
Reducing overfitting. One key difference is the degree of overfitting or the extent to which a model memorizes its training data. Yeom et al. (Yeom et al., 2018) found that defending against membership inference attacks and reducing overfitting go hand-in-hand, so that the goals of privacy and performance are aligned. Many defenses (Shokri et al., 2017; Song and Mittal, 2021; Jia et al., 2019) therefore focus on a regularization effect, and existing regularization methods are also being used to reduce membership risk (Kaya et al., 2020).
  • $\ell^{p}$-regularization and dropout are two of the most commonly used regularization techniques in practice, shown effective against membership inference attacks (Shokri et al., 2017).
  • Early stopping is a straightforward way to prevent overfitting (Song and Mittal, 2021). One halts training when the model’s performance stops improving. This has the effect of stopping the model before it can be begin to memorize.
  • MMD with Mix-up (Li et al., 2021) combines two different mechanisms: (1) a maximum mean discrepancy (MMD) loss to reduce the training accuracy to match validation accuracy, and (2) training on mix-up augmented data, where a linear interpolation between two randomly drawn examples (and their labels) is used instead of individual samples. Mix-up has been shown to increase overall accuracy and further reduce overfitting.
  • MemGuard (Jia et al., 2019) uses an “attack-as-a-defense” approach: It adds carefully-calibrated noise to each confidence score vector predicted by the target model. Once added, the resulting adversarial examples mislead the attacker’s membership identification (e.g., the attack classifiers).
Differential privacy (Dwork, 2006) (DP) is a mathematical framework that provides a probabilistic privacy guarantee. Precisely, the following definition implies that, as a random variable, the privacy loss is bounded in the worst case by a constant $\varepsilon>0$ with probability $10<1-\delta<$ 1.
Definition 2.1 (($\varepsilon,\delta$)-Differential Privacy).
@@
A primary advantage of DP is its immunity to post-processing: an $(\varepsilon,\delta)$-DP algorithm cannot be weakened through manipulation of its output. Based on this property, Abadi et al. proposed an effective, differentially-private adaptation of stochastic gradient descent: DP-SGD (Abadi et al., 2016). The key in DP-SGD is the moment accountant, a mechanism that tracks total privacy leakage possible under a worst-case adversary. This leakage is typically denoted as $\varepsilon$, a convention that will be followed throughout the rest of the paper. Because of the formal privacy guarantee that DP-SGD offers, it has become a de-facto standard for constructing private models. However, this approach inherently leads to a utility loss due to the addition of Gaussian noise to the gradients required for accurate learning.
We evaluate how effective those defenses are in mitigating the membership risks of NODEs and formally demonstrate that NSDEs are DP learners, offering an improved privacy-utility tradeoff.
3. Membership Risks of NODEs
@@
Table 1. Comparison of membership inference risks. We evaluate two different types of neural networks against five existing attacks on CIFAR-10 and CIFAR-100. We compare their risks based on four different metrics used in the prior work.
@@
Acc. (C-10) Acc. (C-100) TPR @ 0.1% FPR TPR @ 1% FPR AUC Inference acc. Model Method Train Test Train Test C-10 C-100 C-10 C-100 C-10 C-100 C-10 C-100 ResNet-14 Yeom et al. (Yeom et al., 2018) 98.1% 86.0% 94.99% 54.26% 0.00% 0.00% 0.00% 0.69% 0.590 0.767 58.92% 73.81% Shokri et al. (Shokri et al., 2017) 0.01% 0.01% 0.14% 0.06% 0.411 0.336 50.00% 50.00% Song and Mittal (Song et al., 2021) 0.00% 0.25% 0.00% 1.76% 0.590 0.494 58.92% 51.19% Watson et al. (Watson et al., 2022) 0.00% 0.00% 0.00% 0.62% 0.590 0.767 58.92% 73.81% Carlini et al. (Carlini et al., 2022) 3.96% 10.52% 10.57% 25.95% 0.679 0.865 61.15% 76.46% NODE (ODENet-16-32-64) Yeom et al. (Yeom et al., 2018) 94.3% 84.4% 82.05% 52.16% 0.00% 0.01% 0.00% 0.71% 0.558 0.701 56.59% 70.75% Shokri et al. (Shokri et al., 2017) 0.04% 0.04% 0.43% 0.37% 0.448 0.398 50.01% 50.00% Song and Mittal (Song et al., 2021) 0.00% 0.15% 0.00% 1.66% 0.529 0.498 54.22% 52.48% Watson et al. (Watson et al., 2022) 0.00% 0.01% 0.00% 0.71% 0.558 0.701 56.69% 70.75% Carlini et al. (Carlini et al., 2022) 1.01% 3.30% 4.22% 12.11% 0.616 0.782 57.65% 72.88%
@@
We first evaluate the membership risks of NODEs. Our hypothesis is that these models are vulnerable to existing membership inference attacks as their performance on popular benchmarking tasks is comparable to conventional feedfoward networks. However, because of their expressivity limited to modeling a system of ODEs, the risk could be lower than that of conventional networks.
3.1. Threat Model
@@
The attacker aims to determine whether a specific sample $z=(x,y)$ is included in the data used for training the target model $f$. The adversary only has black-box access to $f$ [^1] Nasr et al. (Nasr et al., 2019) showed that a white-box membership adversary with full access to the model and its parameters, does not particularly perform better than black-box attacks. We assume that the attacker can train shadow models $f_{s}$ with known membership for all shadow training samples and use them to infer membership from the model output. Because the attacker aims for $f_{s}$ to closely resemble the behaviors of $f$, we assume the attacker’s shadow training dataset and the original training data come from the same underlying data distribution and may overlap partially. We also assume a worst-case black-box adversary with full knowledge of $f$’s architecture and training configurations such that they employ this knowledge to construct and train $f_{s}$. When evaluating defenses in Sec 5, we consider an adaptive adversary who knows the defense mechanisms deployed on $f$ and their hyper-parameters.
3.2. Experimental Setup
@@
Datasets. We use CIFAR-10 and CIFAR-100 (Krizhevsky, 2009), which are object recognition benchmarks, composed of 32$\times$32-pixel RGB images over 10 and 100 object classes, respectively. These datasets are commonly employed in prior work (Yeom et al., 2018; Shokri et al., 2017; Song et al., 2021; Carlini et al., 2022). We employ them to evaluate and compare the membership risks of NODEs with those reported in prior studies on feedforward neural networks.
@@
Models. For each dataset, we train models with architectures based on one studied in Oganesyan et al. (Oganesyan et al., 2020). They have a first few convolutional layer for downsampling, followed by three residual or ODE blocks (see Appendix A for details). For our baseline, we use a ResNet architecture (ResNet-14) with 3 residual blocks where each block within a group has the same size and each group has 16, 32, and 64 filters respectively, widened by a factor of 2. Our NODEs substitute an ODE block for each group of residual blocks. ODENet-16-32-64 has the same dimensions. We also examine ODENet-64, which contains only one ODE block after the downsampling layers—a typical structure used in prior work (Chen et al., 2018; Dupont et al., 2019a).
@@
Metrics. Following best practices from prior work (Carlini et al., 2022), we measure the membership risks by computing the true-positive rate at a low false-positive rate (TPR@0.1–1% FPR). We also compare AUC and inference accuracy (or accuracy). Because this empirical membership risk is highly connected to a model’s generalization gap, we report both the best train and test accuracies of target models.
3.3. Quantifying the Membership Risks
@@
Methodology. We evaluate our models against five membership inference attacks outlined in Sec 2.2. For the attacks employing shadow models, we train 16 models on 16 different subsets of the original training dataset. Following the strategy of prior work (Carlini et al., 2022), we construct these 16 different shadow training datasets such that each training sample is included in half of the sets and excluded from the other half. We then choose one of them as the target model $f$ and the remaining 15 as shadow models $f_{s}$. We run 16-fold cross-validation by choosing a different target model from the 16 models in each round, and we report the averaged metrics. For attacks like those by Yeom et al., which do not use shadow models.
Results. Table 1 summarizes our results. Due to the page limit, we include the ODENet-64 results in Appendix C and compare ResNet-14 and ODENet-16-32-64. We first observe that NODEs exhibit reduced overfitting. To quantify overfitting, we compare the difference between training and testing accuracy. To achieve a comparable test accuracy 84–86% on CIFAR-10, NODEs achieve a training accuracy of 94.3% ($+$9.1%), while ResNet-14 attains 98.1% ($+$12.1%). On a more complex task like CIFAR-100, ResNet-14 shows a difference of 24.6%, compared to 15.8% for NODEs.
Next, we compare the membership risks of NODEs and ResNet-14 models. We observe in overall, NODEs is twice less vulnerable to membership inference attacks, compared to ResNet14, on both CIFAR-10 and CIFAR-100. In most cases on CIFAR-10, attacks other than that by Carlini et al. show a TPR of $\sim$0.0at 0.1% or 1.0% FPR. The attack by Carlini et al. (LiRA) is 2.5–4$\times$ times less effective on NODEs than ResNet-14. On CIFAR-100, the other attacks also show some effectiveness, with TPRs of up to 1.5%, but the LiRA attack remains the most effective. LiRA shows similar effectiveness in at 0.1% FPR, but 1.4$\times$ less effective at 1% FPR. We make the same observations from AUC and inference accuracy.
However, our results do not imply that NODEs are free from membership risks. The attacks are still effective, with TPRs ranging from 1.1% to 10.6% at low FPRs across the two datasets.
@@
Generalization gap is a known factor contributing to membership risks. Previous studies (Yeom et al., 2018; Shokri et al., 2017) indicate that risks increase when a model overfits—it shows a large disparity between the training and testing accuracy. Hence, we analyze whether our results of the reduced membership risks of NODEs is because of overfitting.
Figure 1. Membership risks and overfitting in CIFAR-10.
@@
Figure 1 shows the interaction between membership risks and the generalization gap in CIFAR-10. We plot the attack success and the generalization gap observed for each of the 16 target models, considering LiRA (Carlini et al., 2022), the strongest known attack. Overall, we observe that overfitting has no significant impact on the success of membership inference attacks. In ResNets, the TPR @ 0.1% FPR fluctuates between 3–5% while the generalization gap remains $\sim$12%. NODE models (ODENet-16_32_64 and -64), with 0.5–1.5% TPR @ 0.1% FPR, show the generalization gap in 8–12%. Please refer to Appendix C for our results in CIFAR-100.
3.4. Characterizing the Membership Risks

### v2

@@ 2.2 Membership Inference Attacks
  • Yeom et al. [3] focus on differences in loss values: for a specific instance $z=(x,y)$, a model will be more accurate in its predictions when it has seen $z$ during training. The attack predicts $z$ as a member if and only if, for a threshold $\tau$, the model’s loss on $z$ is below $\tau$; otherwise, it is classified as a non-member.
  • Shokri et al. [4] leverage shadow models which better approximate the differences in a model’s responses to members versus non-members. When constructing shadow models, the attacker artificially generates datasets for training and testing so that the members and non-members are known in advance. The attacker collects responses from shadow models for both members and non-members and trains a classifier to predict membership of $z$ based on the target model’s response.
  • Song and Mittal [5] leverage the prediction correctness to compute the threshold $\tau$ for identifying membership. Instead of training classifiers to perform attacks, $\tau$ is designed such that correct predictions with high confidence yield the lowest values while confident but incorrect predictions achieve the highest.
  • Watson et al. [34] proposed per-example difficulty calibration in which an attacker leverages shadow models trained without a particular example to compute the average confidence level of a model. The average is then subtracted from the example’s confidence obtained from the target model to calibrate the difficulty of the sample.
  • Carlini et al. [6] introduced Likelihood Ratio Attack (LiRA), which refines how shadow training data is used by the adversary. For each sample in the shadow training data, LiRA trains two sets of shadow models: one set with the sample included (“in”) and another without (“out”). The attacker collects logits from both members and non-members across these shadow models, computes a membership score for a sample $z$ as the ratio of the likelihoods of observing its logit in “in” versus “out” models, and uses this score to identify membership.
  • The latest work by Zarifzade et al. [35] introduced a Robust Membership Inference Attack (RMIA), which models membership inference as a fine-grained likelihood-ratio test. RMIA operates by contrasting the target model’s output distribution against a reference distribution, constructed from a small set of reference models and population data. This approach enables more robust differentiation between members and non-members.
Metrics. Initial work [3] uses inference accuracy as a metric for evaluating attack success. Subsequent works use the area under the ROC curve (AUC) as an additional metric to report the balance between true-positive (TPR) and false-positive rates (FPR). More recently, Carlini [6] proposes evaluating attack success in the worst-case low-FPR regime. Our work reports all these metrics in evaluation.
2.3 Defenses against Membership Inference
@@
Reducing overfitting. One key difference is the degree of overfitting or the extent to which a model memorizes its training data. Yeom et al. [3] found that defending against membership inference attacks and reducing overfitting go hand-in-hand, so that the goals of privacy and performance are aligned. Many defenses [4, 5, 36] therefore focus on a regularization effect, and existing regularization methods are also being used to reduce membership risk [24].
  • $\ell^{p}$-regularization and dropout are two of the most commonly used regularization techniques in practice, shown effective against membership inference attacks [4].
  • Early stopping is a straightforward way to prevent overfitting [5]. One halts training when the model’s performance stops improving. This has the effect of stopping the model before it can be begin to memorize.
  • MMD with Mix-up [37] combines two different mechanisms: (1) a maximum mean discrepancy (MMD) loss to reduce the training accuracy to match validation accuracy, and (2) training on mix-up augmented data, where a linear interpolation between two randomly drawn examples (and their labels) is used instead of individual samples. Mix-up has been shown to increase overall accuracy and further reduce overfitting.
  • MemGuard [36] uses an “attack-as-a-defense” approach: It adds carefully-calibrated noise to each confidence score vector predicted by the target model. Once added, the resulting adversarial examples mislead the attacker’s membership identification (e.g., the attack classifiers).
Differential privacy [33] (DP) is a mathematical framework that provides a probabilistic privacy guarantee. Precisely, the following definition implies that, as a random variable, the privacy loss is bounded in the worst case by a constant $\varepsilon>0$ with probability $10<1-\delta<$ 1.
Definition 2.1 (($\varepsilon,\delta$)-DP).
@@
A primary advantage of DP is its immunity to post-processing: an $(\varepsilon,\delta)$-DP algorithm cannot be weakened through manipulation of its output. Based on this property, Abadi et al. proposed an effective, differentially-private adaptation of stochastic gradient descent: DP-SGD [23]. The key in DP-SGD is the moment accountant, a mechanism that tracks total privacy leakage possible under a worst-case adversary. This leakage is typically denoted as $\varepsilon$, a convention that will be followed throughout the rest of the paper. Because of the formal privacy guarantee that DP-SGD offers, it has become a de-facto standard for constructing private models. However, this approach inherently leads to a utility loss due to the addition of Gaussian noise to the gradients required for accurate learning. We evaluate how effective those defenses are in mitigating the membership risks of NODEs and formally demonstrate that NSDEs are DP learners, offering an improved privacy-utility tradeoff.
3 Membership Risks of NODEs
@@
TABLE I: Comparison of membership inference risks. We evaluate two different types of neural networks against six existing attacks on four benchmarks. We compare their risks based on four different metrics used in the prior work.
@@
TPR @ 0.1% FPR TPR @ 1% FPR AUC Inference acc. Model Method F-M C-10 C-100 T-I F-M C-10 C-100 T-I F-M C-10 C-100 T-I F-M C-10 C-100 T-I ResNet-14 Yeom et al. [3] 0.00% 0.00% 0.00% 0.02% 0.00% 0.00% 0.69% 1.11% 0.563 0.590 0.767 0.683 59.91% 58.92% 73.81% 64.27% Shokri et al. [4] 0.05% 0.01% 0.01% 0.10% 0.50% 0.14% 0.06% 1.01% 0.468 0.411 0.336 0.501 50.00% 50.00% 50.00% 50.42% Song and Mittal [38] 0.00% 0.00% 0.25% 0.13% 0.23% 0.00% 1.76% 1.23% 0.471 0.590 0.494 0.499 51.29% 58.92% 51.19% 50.28% Watson et al. [34] 0.00% 0.00% 0.00% 0.12% 0.00% 0.00% 0.62% 1.21% 0.554 0.590 0.767 0.548 56.21% 58.92% 73.81% 53.43% Carlini et al. [6] 1.92% 3.96% 10.52% 0.28% 5.87% 10.57% 25.95% 1.97% 0.599 0.679 0.865 0.554 55.70% 61.15% 76.46% 53.72% Zarifzadeh et al. [35] 1.13% 3.73% 13.98% 0.32% 6.54% 8.78% 28.25% 2.25% 0.497 0.581 0.858 0.578 56.83% 59.13% 75.53% 55.39% NODE Yeom et al. [3] 0.00% 0.00% 0.01% 0.09% 0.06% 0.00% 0.71% 1.12% 0.518 0.558 0.701 0.645 52.55% 56.59% 70.75% 61.23% Shokri et al. [4] 0.09% 0.04% 0.04% 0.10% 0.90% 0.43% 0.37% 1.01% 0.493 0.448 0.398 0.502 50.15% 50.01% 50.00% 50.21% Song and Mittal [38] 0.00% 0.00% 0.15% 0.12% 0.53% 0.00% 1.66% 1.19% 0.507 0.529 0.498 0.500 50.98% 54.22% 52.48% 50.45% Watson et al. [34] 0.00% 0.00% 0.01% 0.12% 0.04% 0.00% 0.71% 1.17% 0.517 0.558 0.701 0.542 52.04% 56.69% 70.75% 53.24% Carlini et al. [6] 0.19% 1.01% 3.30% 0.24% 1.48% 4.22% 12.11% 1.74% 0.529 0.616 0.782 0.543 52.81% 57.65% 72.88% 52.95% Zarifzadeh et al. [35] 0.67% 1.93% 5.41% 0.26% 3.16% 6.57% 17.58% 2.04% 0.523 0.592 0.797 0.570 53.05% 58.17% 73.37% 55.04%
@@
We first evaluate the membership risks of NODEs. Our hypothesis is that these models are vulnerable to existing membership inference attacks as their performance on popular benchmarking tasks is comparable to conventional feedfoward networks. However, because their expressivity is limited to modeling a system of ODEs, the risk could be lower than that of conventional networks.
3.1 Threat Model
@@
The attacker aims to determine whether a specific sample $z=(x,y)$ is included in the data used for training the target model $f$. The adversary only has black-box access to $f$ [^1] Nasr et al. [39] showed that a white-box membership adversary with full access to the model and its parameters, does not particularly perform better than black-box attacks. We assume that the attacker can train shadow models $f_{s}$ with known membership for all shadow training samples and use them to infer membership from the model output. Because the attacker aims for $f_{s}$ to closely resemble the behaviors of $f$, we assume the attacker’s shadow training dataset and the original training data come from the same underlying data distribution and may overlap partially. We also assume a worst-case black-box adversary with full knowledge of $f$’s architecture and training configurations such that they employ this knowledge to construct and train $f_{s}$. When evaluating defenses in §5, we consider an adaptive adversary who knows the defense mechanisms deployed on $f$ and their hyper-parameters.
3.2 Experimental Setup
@@
Datasets. We use four object recognition benchmarks: FashionMNIST [40], CIFAR-10 and CIFAR-100 [41], and Tiny-ImageNet [42]. These datasets are commonly employed in prior work [3, 4, 38, 6]. We employ them to evaluate and compare the membership risks of NODEs with those reported in prior studies on feedforward neural networks.
@@
Models. For each dataset, we train models with architectures based on those studied by Oganesyan et al. [43]. They have a first few convolutional layer for downsampling, followed by three residual or ODE blocks (see Appendix for details). Our baseline is a ResNet architecture (ResNet-14) with 3 residual blocks where each block within a group has the same size and each group has 16, 32, and 64 filters respectively, widened by a factor of 2. To construct our NODE models, we substitute each group of residual blocks with an ODE block. NODE (ODENet-16-32-64) uses the same dimensions. We also examine ODENet-64, which contains only one ODE block after the downsampling layers—a typical structure used in prior work [8, 16].
@@
Metrics. Following best practices from prior work [6], we measure the membership risks by computing the true-positive rate at a low false-positive rate (TPR@0.1–1% FPR). We also compare AUC and inference accuracy (or accuracy). Because this empirical membership risk is highly connected to a model’s generalization gap, we report both the best train and test accuracies of target models.
3.3 Quantifying the Membership Risks
@@
We evaluate our models against six membership inference attacks outlined in §2.2. For the attacks employing shadow models, we train 16 models on 16 different subsets of the original training dataset. Following the strategy of prior work [6], we construct these 16 different shadow training datasets such that each training sample is included in half of the sets and excluded from the other half. We then choose one of them as the target model $f$ and the remaining 15 as shadow models $f_{s}$. We run 16-fold cross-validation by choosing a different target model from the 16 models in each round, and we report the averaged metrics. For attacks like those by Yeom et al., which do not use shadow models.
Results. Table I summarizes our results. Due to the page limit, we include the ODENet-64 results in Appendix and focus on ResNet-14 and NODE models. We contrast the membership risks of NODEs and ResNet-14 models. Overall, we find that NODEs are 2$\times$ less vulnerable to membership inference attacks, compared to ResNet14 in most cases. In CIFAR-10, attacks except those by Carlini et al. and Zarifzadeh et al. show a TPR of $\sim$0.0at 0.1–1.0% FPR. Against the two effective attacks, NODEs reduce their effectiveness by 1.3–4$\times$ than ResNet-14. On CIFAR-100, other attacks also show some effectiveness, with TPRs of up to 1.5%, but the LiRA and RMIA attacks remain the most effective. Both attacks are 1.6–2.6$\times$ less effective on NODEs at the same low-FPR regime. On FashionMNIST and Tiny-ImageNet, we also observe a reduction in membership inference success. However, as the baseline attacks were already less effective on these datasets, the relative reduction is limited at most 2$\times$. We find the reduced effectiveness in AUC and inference accuracy. Note that our results do not imply that NODEs are free from membership risks. The attacks are still effective, with TPRs ranging from 1.1% to 10.6% at low FPRs in CIFAR-10 and 100.
@@
Generalization gap is a known factor contributing to membership risks. Prior work [3, 4] shows that risks increase when a model overfits—it shows a large disparity between the training and testing accuracy. Hence, we analyze whether the reduced membership risks observed in NODEs are because of lower overfitting. Table II summarizes the gaps.
TABLE II: Generalization gap, measured for both ResNet-14 and NODE models across four benchmark datasets.
Task F-M C-10 C-100 T-I Model R-14 NODE R-14 NODE R-14 NODE R-14 NODE Train 100.0% 95.5% 98.1% 94.3% 95.0% 82.1% 64.7% 55.9% Test 88.4% 90.0% 86.0% 84.4% 54.3% 52.2% 42.4% 40.5% $\Delta$ 11.6% 5.5% 12.1% 9.9% 40.7% 29.9% 22.3% 15.4%
Figure 1: Membership risks and overfitting in CIFAR-10.
@@
Figure 1 analyzes the interaction between membership risks and the generalization gap in CIFAR-10. We plot the attack success and the generalization gap observed for each of the 16 target models, focusing on LiRA [6], where we observe the largest reduction. We first find that overfitting correlates with the success of membership inference attacks, as shown in Table II. In ResNets, the generalization gap is $>$12%, and the TPR @ 0.1% FPR is 3–5%. In contrast, NODE models (ODENet-16_32_64 and -64), with 0.5–1.5% TPR @ 0.1% FPR, show the generalization gap in 8–12%. Please refer to Appendix for results in CIFAR-100.
3.4 Characterizing the Membership Risks